OpenStudy (anonymous):

@midhun.madhu1987 Please Help!

OpenStudy (anonymous):

Create a system of equations that includes one linear equation and one quadratic equation. f(x) = 2x + 1 y = x2 + x - 2 Part 1. Show all work to solving your system of equations algebraically. Part 2. Graph your system of equations, and show the solution graphically to verify your solution. I created the two equations I believe I can use, but I am unsure of how to do the two other requirements.

OpenStudy (midhun.madhu1987):

I think this is the way...

OpenStudy (midhun.madhu1987):

The general form of a Linear Equation is: ax + by + c = 0 and that of a Quadratic Equation is: ax^2 + bx + c = 0 Example: x + y = 2(Linear Equation) ---------- (1) x^2 + 2x + 3 = 0 (Quadratic Equation) --------(2) From (1), x = 2 - y Substitute this value of x in equation (2) and solve for y.

OpenStudy (anonymous):

Thanks. Sorry I walked away from the computer for a minute.

OpenStudy (midhun.madhu1987):

Thats okay .. :)

OpenStudy (anonymous):

So what would be an example of a linear equation. I saw the variable version of it you wrote (ax+by+c=0) but what is a numerical version of this?

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (midhun.madhu1987):

i also gave examples also.. equation (1) is a linear equation

OpenStudy (anonymous):

Oh okay I didn't see that. My apologies. Would those be suitable equations to use for this question?

OpenStudy (midhun.madhu1987):

i guess so... u can just check it out

OpenStudy (anonymous):

@midhun.madhu1987 Can you tell me what the ordered pair would be for these solved? x+y=2 x^2+2x+3=0